### v1

# Where to Begin: Efficient Pretraining via Subnetwork Selection and Distillation

Abstract
Small Language models (SLMs) offer an efficient and accessible alternative to Large Language Models (LLMs), delivering strong performance while using far fewer resources. We introduce a simple and effective framework for pretraining SLMs that brings together three complementary ideas. First, we identify structurally sparse sub-network initializations that consistently outperform randomly initialized models of similar size under the same compute budget. Second, we use evolutionary search to automatically discover high-quality sub-network initializations, providing better starting points for pretraining. Third, we apply knowledge distillation from larger teacher models to speed up training and improve generalization. Together, these components make SLM pretraining substantially more efficient: our best model, discovered using evolutionary search and initialized with LLM weights, matches the validation perplexity of a comparable Pythia SLM while requiring $\mathbf{9.2\times}$ fewer pretraining tokens. We release all code and models at https://github.com/whittle-org/whittle/, offering a practical and reproducible path toward cost-efficient small language model development at scale.
1 Introduction
Large Language Models (LLMs) have recently delivered state-of-the-art performance across a wide range of tasks. Their success is largely driven by scale: modern LLMs routinely exceed tens and hundreds of billions of parameters, unlocking remarkable generalization and emergent abilities. However, this scale comes at a cost. Training and deploying such massive models requires substantial computational resources, and inference often exceeds practical memory or latency budgets.
These challenges have motivated increasing interest in Small Language Models (SLMs) (Allal et al., 2025; Yang et al., 2025), which aim to preserve strong performance while remaining deployable in resource-constrained settings such as mobile or edge devices. Although pretraining SLMs is substantially cheaper than training LLMs, the costs are still formidable and often beyond the reach of most smaller research groups. For example, Allal et al. (2025) estimate that training SmolLM2 with 1.7B parameters required on the order of $10^{23}$ FLOPs—roughly \$250,000 of GPU compute.
A common strategy to reduce pretraining cost is to leverage open-weight LLMs as teachers. For instance, Team et al. (2025) used knowledge distillation to train the Gemma 3 family. This idea can be pushed further by warm-starting students from non-random initializations derived from their teachers. Muralidharan et al. (2024) demonstrated this by pruning a teacher model and refining it through distillation, while the smaller variants of Llama 3.2 (Meta AI, 2024) were similarly obtained using a combination of pruning and distillation.
Unfortunately, most existing efforts in this space are closed-source, making them difficult to reproduce and extend. While the evidence so far suggests that teacher models can greatly improve the efficiency of SLM pretraining, the underlying mechanisms remain poorly understood. In this work, we present the first systematic open-source study of warm-starting student models from larger teachers for pretraining. Our contributions are:
[…]
  • Comprehensive analysis. We provide the first systematic comparison of sub-network initialization under knowledge distillation versus standard cross-entropy training. Our study spans multiple student scales and investigates how teacher size influences effectiveness.
  • Reproducible framework. We release an open-source library [^1] https://github.com/whittle-org/whittle/ for extracting sub-networks from existing LLM checkpoints. Together with our empirical findings, this establishes practical guidelines for compute-optimal SLM pretraining across different scales.
Section 2 presents our methodology for extracting sub-networks from a pretrained teacher network, and Section 3 introduces our open-source library for sub-network extraction. We provide an empirical analysis and compare to baseline approaches in Section 4. In Appendix A, we discuss prior work relevant to our approach.
Figure 1: Left: Initialization schemes — random weights, sub-network from a pretrained teacher, and our evolutionary search–based sub-network. Right: The same teacher is used for knowledge distillation to train the student.
Figure 2: Overview of our search spaces and search strategy
2 Methodology
We study the problem of pretraining a Small Language Model (SLM) with the help of a larger open-weight teacher. Our approach follows a two-step strategy: (i) extract a sub-network from the pretrained teacher, and (ii) use this sub-network as initialization for SLM pretraining with knowledge distillation. In this section, we describe the key components of this pipeline. We first introduce the search space granularities considered (Section 2.1), then present our constrained evolutionary search procedure (Section 2.2), and finally delineate the pretraining and distillation process (Section 2.3).
2.1 Search Spaces
We consider a dense transformer model $T$, with $L$ layers and embedding dimension $E$. Each layer $i\in{1,\dots,L}$ consists of a causal self-attention block with $H$ attention heads of dimension $H_{s}$, followed by an MLP block with intermediate dimension $D$. For simplicity, we restrict our discussion to the multi-head attention setting, though the approach extends naturally to multi-query and group-query attention.
We parameterize a sub-network $S$ of the teacher model $T$ by specifying the number of layers $l\in{1,\ldots,L}$, the embedding dimension $e\in{1,\ldots,E}$, the number of attention heads $h\in{1,\ldots,H}$, the head dimension $h_{s}\in{1,\ldots,H_{s}}$, and the MLP intermediate size $d\in{1,\ldots,D}$. We define four search spaces that differ in how weights are selected: coarse versus fine-grained, and uniform versus layer-wise.
Coarse. To construct the sub-network, we always select the first $n$ entries from the corresponding components of the teacher $T$. For example, selecting $h$ attention heads corresponds to taking the first $h$ heads out of the $H$ heads in $T$. Likewise, choosing a smaller embedding dimension $e$ corresponds to taking the first $e$ elements of the embedding vector.
Fine-grained. We select a subset of size $n$ by sampling indices from the teacher’s components. For instance, selecting $h$ attention heads corresponds to sampling $h$ distinct heads from the $H$ available in $T$ (without replacement).
Next, we distinguish between two types of layer configurations:
Uniform. The same configuration $(h,h_{s},d)$ for heads, head size and intermediate MLP size, is applied across all layers. That is, every layer uses the same number of heads, query groups, head dimension, and MLP intermediate size.
Layer-wise. Each layer is allowed to have its own configuration, relaxing the uniformity constraint.
Combining the two sampling strategies (coarse vs. fine-grained) with the two configuration schemes (uniform vs. layer-wise) yields four distinct search spaces:
Coarse Uniform.
This is the simplest search space, in which the same configuration is applied to all layers, always selecting the first entries. For multi-head attention layers, the total number of possible configurations is $N=L\cdot E\cdot H\cdot H_{s}\cdot D$. In the case of group-query attention, $N$ additionally accounts for the number of valid combinations of heads $h$ and query groups $q$.
Coarse Layer-wise.
The coarse layer-wise search space applies coarse sampling independently to each layer in the sub-network $S$, allowing each layer to have its own configuration. The total number of configurations is $N=E\cdot(H\cdot H_{s}\cdot D)^{L},$ which grows exponentially with the number of layers $L$. Compared to the coarse uniform space, which is linear in $L$, the coarse layer-wise space is significantly larger, as each layer can independently select its $(h,h_{s},d)$ configuration.
Fine-grained Uniform .
The fine-grained uniform search space applies fine-grained sampling uniformly across all layers. In this setting, the sub-network may be formed from an arbitrary subset of elements within each layer, rather than being restricted to the first $l$ layers. The total number of configurations in this search space is $N=2^{E\cdot H\cdot H_{s}\cdot D\cdot L}$.
Fine-grained Layer-wise.
The layer-wise fine-grained search space applies fine-grained sampling independently to each layer, yielding the most granular search space considered. Each layer can independently select its number of heads, query groups, head dimension, and MLP intermediate size, and the sub-network may include an arbitrary subset of layers. The total number of configurations is
$N=2^{E}\cdot(2^{H}\cdot 2^{H_{s}}\cdot 2^{D})^{L}\cdot 2^{L},$
which grows exponentially with both the width $(E,H,H_{s},D)$ and the depth $L$, making it the largest and most expressive search space among the variants considered.
2.2 Evolutionary Search
Before outlining our search procedure, we first formalize our experimental setup.
Setup.
Let $\mathcal{M}$ denote a large language model (LLM) parameterized by $\bm{\theta}$, with total parameter count $|\bm{\theta}|=S$ (in billions). We assume $S>1$ and typically consider models where $S>2$. The user specifies a parameter bin
$\mathcal{B}=[S_{\min},S_{\max}],$
which defines the range of acceptable model sizes (e.g., $S_{\min}=1$B, $S_{\max}=2$B).
Parameter-space partitioning.
We partition the overall parameter space
$\mathcal{S}=\{\bm{\theta}:|\bm{\theta}|\in[S_{\min}^{(i)},S_{\max}^{(i)}]\}_{i=1}^{K}$
into $K$ disjoint bins $\{\mathcal{B}_{1},\dots,\mathcal{B}_{K}\}$, each corresponding to a contiguous range of parameter counts. This stratification ensures balanced coverage across different model sizes. Without such binning, uniform random sampling tends to under-represent very small and very large models.
We now delineate our constrained evolutionary search procedure.
Evolutionary search with constraint enforcement.
Within each bin $\mathcal{B}_{i}$, we perform an evolutionary search over candidate sub-network architectures $\mathcal{A}(\bm{\theta})$. At each iteration, candidate architectures are sampled and evaluated according to a fitness function $f(\mathcal{A})$. To enforce the bin constraint, we apply rejection sampling:
$\mathcal{A}\leftarrow\begin{cases}\mathcal{A},&\text{if }|\bm{\theta}_{\mathcal{A}}|\in\mathcal{B}_{i},\\[3.0pt] \text{reject},&\text{otherwise}\end{cases}$
Only candidates satisfying $|\bm{\theta}_{\mathcal{A}}|\in\mathcal{B}_{i}$ are retained for further evolution.
Selection.
After convergence, we return the set of small language models (SLMs)
$\mathcal{A}^{*}_{i}=\arg\max_{|\bm{\theta}_{\mathcal{A}}|\in\mathcal{B}_{i}}f(\mathcal{A}),$
that achieve the most favorable initialization for subsequent pretraining or distillation. These selected SLMs represent the optimal sub-network architectures within the specified parameter range.
The overall procedure is summarized in Algorithm 1 (Appendix C.2). Within each parameter-size bin, we initialize a population of sub-network candidates drawn from the constrained search space. At each epoch, candidates are evaluated by perplexity and the top-$k$ elites are retained. Genetic operators—mutation and crossover—then generate offspring subject to bin constraints, while additional random samples encourage exploration. The next population is formed by selecting the $N$ best candidates among elites, offspring, and random samples. After $T$ epochs, the best-performing sub-network in each bin is returned, with mutation and crossover formally defined below.
Mutation.
Given a candidate architecture $\mathcal{A}$, we define a mutation operator $\mu(\mathcal{A})$ that perturbs one architectural dimension at a time. Specifically, we uniformly sample a dimension
$x\in\{l,e,h,g,d,h_{s}\},$
where $l$ denotes the number of layers, $e$ the embedding dimension, $h$ the number of attention heads, $g$ the number of query groups, $d$ the intermediate (feedforward) dimension, and $h_{s}$ the per-head size.
Mutation in layer-wise search space.
In layer-wise search spaces, architectural attributes $(h,g,d,h_{s})$ are defined independently for each layer. A mutation of the layer count $l\to l^{\prime}$ is handled as follows:
[…]
If $x\in\{h,g,d,h_{s}\}$, we first sample a layer index $i\sim\text{Uniform}\{1,\dots,l\}$, then resample the chosen dimension for that layer:
$x_{i}^{\prime}\sim\text{Uniform}\{\text{choices}(x)\}$
Where choices(x)(x) denotes the valid choices for an architectural attribute $x$ defined by a search space. All other architectural parameters remain fixed.
Mutation in Fine-grained search space.
In the fine-grained setting, mutations operate at the neuron level. For instance, mutating the embedding dimension $e\to e^{\prime}$ corresponds to
[…]
This formulation enables smooth exploration of architectures across both coarse (layer-wise) and fine-grained structural variations, while maintaining consistency with the model size constraint.
Crossover.
To produce a child architecture from two parents, $P_{1}$ and $P_{2}$, we apply a crossover operator $\chi(P_{1},P_{2})$. We first require both parents to share the same number of layers:
$l_{P_{1}}=l_{P_{2}}=l$
Let the architectural dimensions of each parent be
$P_{1}=(e_{1},h_{1},g_{1},h_{s,1},d_{1}),\quad P_{2}=(e_{2},h_{2},g_{2},h_{s,2},d_{2}),$
where $e$, $h$, $g$, $h_{s}$, and $d$ denote the embedding dimension, number of attention heads, number of query groups, head size, and intermediate dimension, respectively.
A child architecture $c$ is then generated by independently inheriting each dimension from one of the two parents:
$x_{c}=\begin{cases}x_{1},&\text{with probability }0.5,\\[3.0pt] x_{2},&\text{with probability }0.5,\end{cases}\quad\text{for each }x\in\{e,h,g,h_{s},d\}$
For example, a valid crossover outcome might be
$c=(e_{2},h_{1},g_{2},h_{s,2},d_{1})$
This independent dimension-wise crossover enables fine-grained recombination of architectural traits while preserving structural compatibility (e.g., layer count consistency) between parents.
2.3 SLM Pretraining and Distillation
Sub-network Extraction.
Our constrained evolutionary search Algorithm 1 (Appendix C.2), returns a sub-network configuration $s_{b}$, for every parameter bin $b$. Given this sub-network configuration, we extract the smaller language model corresponding to this configuration from the larger base model we perform search on. We then convert this extracted sub-network into a dense language model with the corresponding architecture. This is then the SLM that we use in our pretaining pipeline, optimizing the standard token-level cross entropy, language modeling loss.
Knowledge Distillation.
Model distillation (Hinton et al., 2015), or knowledge distillation, compresses a large teacher model into a smaller student network that achieves similar performance with fewer resources. Instead of training solely on hard labels, the student leverages soft labels from the teacher, obtained via temperature-scaled softmax:
$p_{i}^{(T)}(z)=\frac{\exp\!\left(\frac{z_{i}}{T}\right)}{\sum_{j}\exp\!\left(\frac{z_{j}}{T}\right)},$ (1)
where $z=[z_{1},z_{2},\dots,z_{n}]$ are logits and $T>0$ is the temperature. The student parameters $\hat{{\mathbf{w}}}_{\theta}$ are optimized with a loss combining hard-label cross-entropy and distillation:
$\mathcal{L}=\alpha\,\mathcal{L}_{\text{CE}}(\mathbf{y},\mathbf{s})+\beta\,\mathcal{L}_{\text{D}}(p_{t},p_{s}),$ (2)
where $\alpha,\beta\in[0,1]$ ,and $p_{t}$ and $p_{s}$ are the teacher and student logit distributions, respectively. In our setting, $\mathcal{L}_{\text{D}}$ is the forward KL divergence,
$\mathcal{L}_{\text{D}}=\sum_{i}p_{t_{i}}^{(T)}\log\frac{p_{t_{i}}^{(T)}}{p_{s_{i}}^{(T)}},$ (3)
encouraging the student distribution $p_{s}^{(T)}$ to match the teacher’s softened distribution $p_{t}^{(T)}$. In our final knowledge-distillation setup, we use equation 2, with $\mathcal{L_{D}}$, corresponding to the forward-kl divergence depicted in equation 3.
Top-$k$ Logit Distillation.
We define a variant of knowledge distillation that truncates the teacher distribution to its $k$ most salient outputs. Let $z_{t},z_{s}\in\mathbb{R}^{C}$ denote the teacher and student logits, respectively, and $T$ the temperature parameter. The teacher distribution is given by $p_{t}=\text{softmax}(z_{t}/T)$. We denote by $\mathcal{K}\subset\{1,\dots,C\}$ either the indices of the top-$k$ logits of $z_{s}$, or a subset sampled from $p_{t}$. The distillation loss is then defined as:
$\mathcal{L}_{\text{top-}k}=\sum_{i\in\mathcal{K}}\text{KL}\!\left(\text{softmax}\!\left(\tfrac{z_{t}^{(i)}}{T}\right)\,\|\,\text{softmax}\!\left(\tfrac{z_{s}^{(i)}}{T}\right)\right).$ (4)
3 Whittle: A Library for SLM Pre-training and Distillation
Recent model releases such as LLaMA 3.1–8B, LLaMA 3.2–1B, and LLaMA 3.2–3B [^2] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ leverage pruning and distillation to produce smaller variants, but their training recipes and code are closed-source, hindering reproducibility. Similarly, Minitron (Muralidharan et al., 2024) outlines best practices for SLM pretraining, but its implementation [^3] https://github.com/NVIDIA-NeMo/NeMo/tree/main is not readily generalizable across model families.
To address this gap, we present whittle, a fully open-source library that provides a reproducible, general-purpose pipeline for extracting and pretraining SLMs directly from Hugging Face models. Whittle supports a range of functionalities to allow for flexible search space design, sub-network search, extraction, pretraining and knowledge distillation. In this section, we outline the core functionalities of whittle and its API design:
set_sub_network().
Given a pretrained decoder-only LLM from litgpt, we first convert it into a whittle model to enable flexible sub-network extraction. To evaluate a sub-network using the whittle model, we dynamically activate only structured components of the LLM associated with that sub-network using the set_sub_network() API (Listing 1). It allows the user to explicitly set architectural parameters of the sub-network, such as embedding dimension, intermediate size, number of heads, layers, query groups, and head size, as well as indices for sampled neurons, layers, and heads. Importantly, it allows to vary the number of heads, head size, intermediate size, and query groups across layers. This function is a core utility in whittle, supporting downstream procedures such as search, pretraining, and distillation.
search().
The search() API (Listing 3, Appendix F) constructs a whittle super-network from a base HuggingFace model and facilitates automated sub-network selection. It supports evolutionary strategies as well as algorithms from syne-tune (Salinas et al., 2022) [^4] https://github.com/syne-tune/syne-tune, and performs constrained search across parameter bins via rejection sampling. Each candidate sub-network is instantiated through set_sub_network() and evaluated on a task-specific metric, such as perplexity, to guide the search process.
convert_subnet_to_litgpt_model().
The convert_subnet_to_litgpt_model() function (Listing 4, Appendix F) transforms a selected sub-network configuration into a standalone GPT model within the litgpt framework. Given a super-network and a dictionary specifying architectural configurations (e.g., embedding dimension and number of heads), this utility extracts the corresponding sub-network and instantiates it as an independent GPT model. The resulting model can then be employed for downstream tasks such as pretraining, fine-tuning, or distillation.
pretrain().
The pretrain() function (Listing 2, Appendix F) enables pretraining of a sub-network initialized from a checkpointed GPT model. Given the model weights, a configuration file describing the sub-network architecture, and a target dataset, this utility restores the model and resumes training from the specified state.
distill().
The distill() function (Listing 5, Appendix F) supports knowledge distillation from a larger teacher model into a sub-network extracted from a checkpoint. Given a teacher model, sub-network configuration, and a target dataset, this utility trains the sub-network under the supervision of a specified teacher (e.g., EleutherAI/pythia-12b). Different distillation objectives (e.g., forward KL divergence) and constraints such as top-$k$ token selection are supported.
4 Experiments
Our study focuses on the Pythia (Biderman et al., 2023) family of models, which span sizes from 14M to 12B parameters; in particular, we use the 6.9B and 12B variants. Importantly, the modular design of our framework ensures that the methodology is readily applicable to any large language model supported by litgpt [^5] https://github.com/Lightning-AI/litgpt/.
Our experiments are organized around three core components: (a) sub-network search, (b) pretraining of small language models (SLMs), and (c) distillation into SLMs. For each component, we outline the setup, present results, and highlight key insights. We now discuss them in turn.
4.1 Search Space Definitions
Table 1 summarizes the search spaces for Pythia-6.9B and Pythia-12B, listing the allowable values for each transformer dimension. In the coarse layer-wise and fine-grained layer-wise settings, $h$, $h_{s}$, and $d$ are sampled independently at each layer. The fine-grained spaces extend this further with neuron-level sampling within each dimension, as detailed in Section 2.1.
Table 1: Search space configurations for different model families. Here, $e$ denotes embedding dimension, $h$ the number of attention heads, $h_{s}$ the head size, $l$ the number of layers and $d$ the MLP dimension
Base Model $e$ $h$ $h_{s}$ $l$ $d$
EleutherAI/pythia-6.9b $[1,4096]$ $[1,32]$ $\{4,6,8,\dots,128\}$ $[1,32]$ $[1,16384]$
EleutherAI/pythia-12b $[1,5120]$ $[1,40]$ $\{4,6,8,\dots,128\}$ $[1,36]$ $[1,20480]$
4.2 Evolutionary Search for Optimal SLMs
Search Setup
We apply Algorithm 1 (Appendix C.2) to conduct evolutionary search over parameter bins in Pythia-6.9B and Pythia-12B, considering the coarse uniform, coarse layer-wise, fine-grained uniform, and fine-grained layer-wise [^6] In the fine-grained spaces, $h$, $h_{s}$, and $d$ are sampled independently at each layer, with additional neuron-level sampling within each dimension. See Section 2.1 for details. search spaces from Section 2.1.
Figure 3: Best perplexity after evolutionary search based on perplexity for different search spaces.
We use perplexity on wikitext (Merity et al., 2017) as selection metric, with mutation and crossover probabilities fixed at $0.2$. For Pythia-6.9B and 12B, we define three bins with parameter counts of 385M–426M (bin-1), 961M–1.06B (bin-2), and 2.64B–2.91B (bin-3), centered at 5% of Pythia-410M, Pythia-1B, and Pythia-2.8B, respectively [^7] The bins were computed based on the exact number of parameters in the Pythia models. Each evolutionary run proceeds for 100 epochs, with results in the fine-grained layer-wise setting reported at the final epoch before rejection sampling becomes infeasible due to the combinatorial growth of the search space.
Results Discussion.
Figure 3 reports the perplexity of pruned sub-networks from Pythia-6.9B across bin-1, bin-2, and bin-3 under different search spaces on the wikitext test set. Note that these sub-networks are evaluated without any further pretraining or finetuning. We observe that searches constrained to the smaller coarse uniform and coarse layer-wise spaces generally yield more effective sub-networks.
4.3 Pretraining of SLMs
Pretraining Setup.
[…]
Results and Discussion.
Figure 4 presents the pretraining results. We compare pretraining of the extracted best sub-network (Supernet-init) against two baselines: (i) Random-init, where the same architecture is trained with random initialization and a 10B-token budget, and (ii) the original Pythia model (center of the bin), also trained with random initialization and the same budget. Across parameter bins, initializing from the supernet yields consistent improvements in validation perplexity. Notably, the gains are most pronounced for bin-3, indicating that supernet initialization is particularly beneficial in higher-parameter regimes, where our model achieved the same validation perplexity with $9.2\times$ fewer total training tokens.
Figure 4: Pretraining Validation perplexity of the best sub-networks from each search space and the bin-center Pythia models (410M, 1B, 2.8B), all trained for 10B tokens with cross-entropy loss. Sub-networks are extracted from the Pythia-6.9B base model.
4.4 Distillation of SLMs
Distillation Setup.
To further accelerate convergence, we distill knowledge from Pythia-6.9B and Pythia-12B teacher models. As described in Section 2, training is performed with a weighted combination of forward-KL divergence and cross-entropy loss ($0.8$ and $0.2$, respectively). For computational efficiency, we apply top-$k$ logits distillation with $k=1024$ and a distillation temperature of $0.9$. For distillation, we select the best architectures from every bin, determined by pretraining for a small token budget of 2B tokens in section4.3, and train it with a larger token budget of 10B tokens with the distillation loss function on Nemotron-CC. When training a sub-network with distillation loss, we use the same model for the teacher as the one that the sub-network was extracted from (a sub-network extracted from Pythia-6.9B uses the Pythia-6.9B as the teacher model as well).
Figure 5: Distillation: Comparison of validation perplexity for models trained with distillation loss v/s cross entropy loss. All sub-networks are extracted from Pythia-6.9B as a base model and trained for 10B tokens
Results and Discussion.
Figure 5 illustrates the effect of distillation. We find that distillation consistently improves perplexity in both bin-1 and bin-2, with the model in bin-2 achieving the same performance as Pythia-1B with $5.11\times$ fewer tokens.
4.5 Evaluation on Downstream Tasks
Base Model Initialization #Params COPA Lambada-OpenAI Winogrande MMLU PIQA ARC-challenge ARC-easy HellaSwag Avg-acc PPL-Nemotron-$c$
Pythia-6.9B Random Init 389M 59.00 18.51 51.14 26.43 63.87 24.74 46.80 31.28 37.77 26.20
Supernet Init 389M 61.00 24.02 51.54 26.35 65.34 24.57 51.38 33.11 39.33 23.66
Pythia-12B Random Init 407M 57.00 14.87 50.67 26.33 61.64 23.97 42.47 29.76 36.06 29.57
Supernet Init 407M 63.00 18.37 52.09 25.99 62.73 23.55 46.42 30.91 37.50 27.33
Pythia-410M* Random Init 405M 62.00 19.54 50.67 25.54 64.14 24.57 47.35 32.70 38.42 25.29
Pythia-6.9B Random Init 1.04B 64.00 23.36 51.54 26.62 65.78 27.13 51.80 36.83 40.22 21.84
Supernet Init 1.04B 66.00 38.52 51.46 26.09 69.26 30.12 63.51 45.20 43.74 17.77
Pythia-12B Random Init 1.04B 63.00 23.33 50.51 26.14 66.76 26.36 53.74 36.65 39.32 21.21
Supernet Init 1.04B 64.00 27.56 51.77 26.19 66.54 26.45 53.96 36.42 40.88 20.77
Pythia-1B* Random Init 1.01B 64.00 25.67 52.41 25.20 66.00 28.24 56.14 38.23 41.06 20.11
Pythia-6.9B Random Init 2.91B 61.00 26.49 52.10 26.39 67.74 28.33 57.83 41.12 41.31 13.75
Supernet Init 2.91B 66.00 50.16 56.91 26.45 72.69 33.87 67.09 53.40 47.41 10.99
Pythia-12B Random Init 2.91B 67.00 27.32 50.36 25.25 67.85 27.65 57.79 40.58 41.73 13.26
Supernet Init 2.91B 69.00 41.76 51.46 26.20 70.57 28.67 61.99 45.98 44.47 11.71
Pythia-2.8B* Random Init 2.78B 68.00 24.51 53.03 25.74 67.68 25.34 46.67 39.11 40.55 14.54
Table 2: Evaluation of sub-networks extracted from Pythia-6.9B and Pythia-12B after pretraining on 10B tokens. A Pythia model of comparable size is also trained on the same budget with random initialization to serve as a baseline (indicated with *). Reported numbers are metrics as defined in Section 4.5 (%).
Evaluation Setup.
We evaluate our pretrained and distilled sub-networks on different commonsense and question-answering type tasks. Specifically, we evaluate 0-shot performance on copa, lambada_openai, and winogrande, 5-shot performance on MMLU, and 10-shot performance on arc-easy, arc_challenge, piqa, and hellaswag. We report accuracy for copa, lambada_openai, winogrande, MMLU and length normalized accuracy for piqa, arc_easy, arc_challenge and hellaswag. We use lm-eval-harness [^8] https://github.com/EleutherAI/lm-evaluation-harness to perform evaluation on downstream tasks.
Results Discussion.
Table 2 reports average downstream accuracies for our best sub-networks in each parameter bin pretrained with a 10B-token budget. For comparison, we also include Pythia-410M, 1B, and 2.8B models trained with the same budget. Across all bins, Supernet-init outperforms both Random-init (for the same extracted architecture) and the original Pythia architectures (bin centers). Furthermore, sub-networks extracted from the smaller base model (Pythia-6.9B) consistently outperform those from the larger base (Pythia-12B). We present results of our distilled models on downstream tasks in Table 10.
5 Ablations
In this section, we conduct ablation studies to examine the effect of four key factors in our framework: (a) the choice of search space, (b) the loss function used for distillation, (c) the performance metric employed during search.
Granularity of Search Spaces.
Figure 6 illustrates the effect of varying search space granularity. We find that different bins benefit from distinct choices: for bin-1, fine-grained uniform search space is optimal; for bin-2, coarse layer-wise performs best; and for bin-3, coarse uniform yields the strongest results.
Figure 6: Validation perplexity of the best models from each search space found via evolutionary search. All models are initialized with Pythia-6.9B weights and trained for 2 billion tokens. Within each bin, the models’ parameter counts fall within a $\pm 5\%$ range of that bin’s target size.
Full logits vs. top-$k$ logits.
In our distillation experiments in Section 4, following (Team et al., 2025), we use top-k logit based distillation. Here, we ablate this choice for the distillation loss by comparing supervision from the full teacher distribution against a truncated variant using only the top-$k$ logits (Figure 8). This isolates how much of the teacher’s probability mass is required for effective transfer. We find that, in general, distilling from the full-logit distribution yields a lower perplexity.
Figure 7: Full vs. top-$k$ logit distillation.
Figure 8: Comparison of search guided by importance metrics and perplexity. We report results in the best search space for each bin.
Metric for Searching sub-networks.
Finally, in Figure 8, we evaluate different search metrics. Specifically, we compare activation-based importance scores (as in Minitron (Muralidharan et al., 2024)) and weight-magnitude scores (Han et al., 2015) against directly optimizing for perplexity in our setup. We define the details of the importance score computation procedure, i.e. the metric guiding the search, in Appendix E. All searches are run with the same 16-hour budget per parameter bin. We find that perplexity-based search consistently achieves lower perplexity than proxy metrics, suggesting that importance and magnitude scores are less reliable indicators of sub-network quality.
6 Conclusion
We present a principled framework for initializing small language models (SLMs) by extracting sub-networks from a larger pre-trained teacher network. Our experiments demonstrate that this approach accelerates the overall pre-training process of SLMs by up to $9.2\times$ compared to baseline SLM models of similar size. To select the sub-network, we employ a constrained evolutionary search strategy that identifies optimal candidates based on validation performance. Further, we analyze four different search spaces of increasing granularity and demonstrate that for the larger variants of SLMs, the least granular search space (coarse uniform) yields the best model. The smaller variants, however, benefit from more granular search spaces such as fine-grained uniform and coarse layer-wise.
For future work, we aim to derive scaling laws to better understand the impact of improved initialization strategies as model and data scales increase. Additionally, we plan to investigate the effect of teacher model choice on student performance, particularly in domain-specific settings. For example, it remains an open question whether a multilingual teacher provides advantages over an English-only teacher when training a monolingual student model.
Acknowledgments
This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828 and 499552394, SFB 1597 (SmallData); the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them. Frank Hutter acknowledges the financial support of the Hector Foundation. The authors gratefully acknowledge the computing time provided on the high-performance computer HoreKa by the National High-Performance Computing Center at KIT (NHR@KIT). This center is jointly supported by the Federal Ministry of Education and Research and the Ministry of Science, Research and the Arts of Baden-Württemberg, as part of the National High-Performance Computing (NHR) joint funding program (https://www.nhr-verein.de/en/our-partners). HoreKa is partly funded by the German Research Foundation (DFG)
Appendix A Related Work
Model Pruning.
Pruning is a core approach for compressing neural networks by removing redundant parameters while preserving accuracy. Early work on unstructured magnitude pruning [LeCun et al., 1990, Han et al., 2016] achieved high sparsity with minimal accuracy loss, but offered limited inference benefits on modern hardware. This motivated structured and semi-structured pruning methods that remove neurons, filters, or enforce hardware-friendly sparsity patterns [Li et al., 2017, Zhou et al., 2021, Ma et al., 2023, Frantar and Alistarh, 2023]. The Lottery Ticket Hypothesis (LTH) [Frankle and Carbin, 2018] provided a compelling rationale, showing that large networks contain sub-networks (“winning tickets”) that can train in isolation to match full-model performance. Subsequent work examined their generalization across architectures and optimizers [Morcos et al., 2019, Desai et al., 2019], their stabilization and theoretical underpinnings [Frankle et al., 2019, Malach et al., 2020], and their presence in large pretrained language models [Chen et al., 2020, Prasanna et al., 2020, Liang et al., 2021]. These advances highlight pruning as a powerful tool for efficient deployment in resource-constrained settings. Central to both pruning and ticket discovery is the design of importance scores—criteria based on weight magnitude, gradients, or activations [Molchanov et al., 2019, Frantar and Alistarh, 2023, An et al., 2024] that estimate which components can be removed with minimal loss. However, efficiently scaling such methods to billion-parameter LMs remains a major challenge. Our work addresses this gap by introducing a framework for discovering high-quality sub-networks that is efficient, scalable, and easily parallelizable.
Knowledge Distillation (KD).
KD compresses large language models by transferring knowledge from a teacher to a smaller student, aiming to preserve accuracy while reducing compute [Hinton et al., 2015, Xu et al., 2024]. For autoregressive LMs, this is typically done in two ways. Logit-based distillation trains the student to match the teacher’s output distribution via KL-divergence, often with top-$k$ or top-$p$ truncation to mitigate noise from heavy-tailed distributions [Hinton et al., 2015, Kim and Rush, 2016, Sanh et al., 2019, Team et al., 2024]. Representation-based distillation instead aligns internal dynamics, training the student to mimic hidden states or their projections using MSE losses [Romero et al., 2015, Jiao et al., 2020, Wang et al., 2020b]. These complementary strategies highlight KD’s versatility in shaping both outputs and internal representations. Beyond compression, KD smooths decision boundaries and provides richer training signals, often yielding faster and more stable convergence. Building on these insights, we demonstrate the effectiveness of KD as a key ingredient for efficient SLM pretraining.
Neural Architecture Search (NAS).
NAS [White et al., 2023, Elsken et al., 2019b] automates the exploration of large architecture spaces. Existing approaches include black-box optimization [Zoph and Le, 2017, White et al., 2021, Real et al., 2019, Shen et al., 2023, Zhou et al., 2019, Schrodi et al., 2023], which repeatedly train and evaluate candidates, and gradient-based methods [Liu et al., 2019, Dong and Yang, 2019, Chen et al., 2021, Zela et al., 2020], which perform differentiable search over a weight-sharing supernetwork. Extensions incorporate hardware-awareness and multi-objective criteria [Sukthanker et al., 2025, Elsken et al., 2019a, Lu et al., 2019, Hsu et al., 2018, Lu et al., 2020, Sukthanker et al., 2024, Lee et al., 2021, Li et al., 2021, Klein et al., 2024], jointly optimizing accuracy, efficiency, and deployment constraints. A major limitation, however, is the expensive supernet pretraining required by most methods [Cai et al., 2020, Sukthanker et al., 2024, Wang et al., 2020a], which is prohibitive at the scale of LLMs. Our approach sidesteps this by leveraging open-source pretrained LLMs as the basis for search, eliminating supernet pretraining. Moreover, unlike traditional NAS that seeks architectures for direct deployment, we focus on discovering sub-networks that provide strong initializations for efficient pretraining.
SLM Pretraining in Practice.
Recent open-source releases often provide families of models ranging from compact Small Language Models (SLMs) to much larger variants. SLMs are especially important for edge deployment, where efficiency and memory are critical. A straightforward way to obtain them is to train models across multiple scales [Biderman et al., 2023], but this is computationally costly. To reduce training demands, recent work instead trains a large base model and extracts smaller ones via pruning and distillation [Muralidharan et al., 2024, Meta AI, 2024, Team et al., 2025], or relies solely on distillation from a larger teacher, as in Gemma-3 [Team et al., 2025]. Despite this progress, there remains no principled framework for compute-efficient SLM pretraining. Our work addresses this gap through a systematic study of sub-network extraction and initialization strategies, combined with pipeline designs and loss functions for training high-performing SLMs.
Appendix B Hyperparameter Configurations of Experiments
In Tables 4 - 8, we present the hyperparameter settings for all our experiments.
Table 3: Hyperparameters used for the Best Performing Subnets per Bin (Parameter Range) from the Pythia-12B Model
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Coarse Bin 1 385M–426M Model & Data Model Name pythia-12b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 8 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Layerwise Bin 2 961M–1.06B Model & Data Model Name pythia-12b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 8 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Bin 3 2.64B–2.91B Model & Data Model Name pythia-12b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $1.6\times 10^{-4}$ Min Learning Rate $1.6\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 16 LR Warmup Steps 238 Max Sequence Length 2048 Seed 42
Table 4: Hyperparameters used for the Best Performing Subnets per Bin (Parameter Range) from the Pythia-6.9B Model
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Model Name pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Layerwise Bin 2 961M–1.06B Model & Data Model Name pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 4 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Bin 3 2.64B–2.91B Model & Data Model Name pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $1.6\times 10^{-4}$ Min Learning Rate $1.6\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Training & Batching Total Training Tokens 50B Global Batch Size 1056 Micro Batch Size 16 LR Warmup Steps 238 Max Sequence Length 2048 Seed 42
Table 5: Hyperparameters used for Distillation Experiments on the Best Performing Subnets per Bin (Parameter Range) from the Pythia-12B Model
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Coarse Bin 1 385M–426M Model & Data Teacher Model pythia-12b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 2 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Layerwise Bin 2 961M–1.06B Model & Data Teacher Model pythia-12b Precision bf16-true Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 8 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Bin 3 2.64B–2.91B Model & Data Teacher Model pythia-12b Precision bf16-true Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $1.6\times 10^{-4}$ Min Learning Rate $1.6\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 4 LR Warmup Steps 238 Max Sequence Length 2048 Seed 42
Table 6: Hyperparameters used for Distillation Experiments on the Best Performing Subnets per Bin (Parameter Range) from the Pythia-6.9B Model
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Layerwise Bin 2 961M–1.06B Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 4 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Coarse Bin 3 2.64B–2.91B Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $1.6\times 10^{-4}$ Min Learning Rate $1.6\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 10B Global Batch Size 1056 Micro Batch Size 4 LR Warmup Steps 238 Max Sequence Length 2048 Seed 42
Table 7: Hyperparameters used for Distillation Ablation Experiments on a Subnet from the Pythia-6.9B Model using varying Teacher Model Sizes
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Teacher Model pythia-1b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 2B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 2B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42
Table 8: Hyperparameters used for Distillation Ablation Experiments on a Subnet from the Pythia-6.9B Model — Top-K vs. Full Logits
Search Space Parameter range Hyperparameter Type Value Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Top-1024 Training & Batching Total Training Tokens 2B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42 Evolutionary Search Finegrained Bin 1 385M–426M Model & Data Teacher Model pythia-6.9b Precision bf16-mixed Dataset Nemotron-CC Optimizer & Regularization Optimizer AdamW Learning Rate $3\times 10^{-4}$ Min Learning Rate $3\times 10^{-5}$ Weight Decay $0.01$ AdamW $\beta_{1},\beta_{2}$ $0.9,0.95$ Gradient Clipping Norm $1.0$ Distillation $\alpha$ $0.2$ $\beta$ $0.8$ Temperature $0.9$ Logits Full Training & Batching Total Training Tokens 2B Global Batch Size 1056 Micro Batch Size 6 LR Warmup Steps 0 Max Sequence Length 2048 Seed 42
Appendix C Additional Methodological Details
C.1 Attention masking
The attention mechanism used in transformer blocks naturally supports sub-network extraction. In practice, this means that an attention mechanism can be masked to yield a smaller, distinct type of attention. Figure 10 provides an overview of the main variants—multi-head attention (MHA), multi-query attention (MQA), and grouped-query attention (GQA). Since GQA serves as a super-class of these mechanisms, it can be transformed into either MHA or MQA. An illustration of this transformation is shown in Figure 11.
C.2 Evolutionary Search Algorithm
We present the details of our evolutionary search algorithm in Algorithm 1.
Figure 9: An overview of the Whittle library.
Figure 10: An illustration of the different types of attention mechanisms. In multi-head attention (MHA), each query is paired with its own key and value; in multi-query attention (MQA), multiple queries share a single key–value pair; and in grouped-query attention (GQA), multiple key–value pairs are used, with each pair serving more than one query.
Figure 11: An example of how grouped-query attention (GQA) can be masked to emulate other forms of attention, such as multi-head or multi-query attention. The masked heads are shown in gray. Note that GQA can also be reduced to fewer query heads while preserving the same number of groups.
Appendix D Additional Results
Below, we present additional experimental results. Figure 12 shows the effect of using different teachers for knowledge distillation, Figure 13 shows the evolutionary search trajectory for different parameter bins, with the best perplexity marked in red. Table 10, presents the result of distilled models on downstream tasks. Table 9, provides the results on an extended set of commonsense reasoning based downstream tasks.
Base Model Initialization #Params COPA OpenBookQA Lambada-OpenAI Winogrande Social IQA MMLU-cont. MMLU CommonsenseQA PIQA ARC-challenge ARC-easy HellaSwag BoolQ Avg-acc PPL-Nemotron-$c$
Pythia-6.9B Random Init 389M 59.00 29.20 18.51 51.14 36.59 25.82 26.43 19.41 63.87 24.74 46.80 31.28 58.17 37.77 26.20
Supernet Init 389M 61.00 30.00 24.02 51.54 38.23 26.21 26.35 18.84 65.34 24.57 51.38 33.11 60.73 39.33 23.66
Pythia-12B Random Init 407M 57.00 27.80 14.87 50.67 35.93 25.32 26.33 20.48 61.64 23.97 42.47 29.76 52.50 36.06 29.57
Supernet Init 407M 63.00 27.40 18.37 52.09 37.10 25.90 25.99 21.21 62.73 23.55 46.42 30.91 52.78 37.50 27.33
Pythia-410M Random Init 405M 62.00 29.60 19.54 50.67 36.49 25.72 25.54 20.15 64.14 24.57 47.35 32.70 61.01 38.42 25.29
Pythia-6.9B Random Init 1.04B 64.00 29.20 23.36 51.54 38.59 26.63 26.62 21.05 65.78 27.13 51.80 36.83 60.36 40.22 21.84
Supernet Init 1.04B 66.00 34.60 38.52 51.46 41.04 28.98 26.09 19.81 69.26 30.12 63.51 45.20 53.98 43.74 17.77
Pythia-12B Random Init 1.04B 63.00 28.40 23.33 50.51 37.92 27.05 26.14 19.57 66.76 26.36 53.74 36.65 51.77 39.32 21.21
Supernet Init 1.04B 64.00 31.20 27.56 51.77 38.48 27.36 26.19 19.82 66.54 26.45 53.96 36.42 61.65 40.88 20.77
Pythia-1B Random Init 1.01B 64.00 30.20 25.67 52.41 38.95 26.93 25.20 20.97 66.00 28.24 56.14 38.23 60.83 41.06 20.11
Pythia-6.9B Random Init 2.91B 61.00 30.60 26.49 52.10 39.00 27.60 26.39 19.82 67.74 28.33 57.83 41.12 59.05 41.31 13.75
Supernet Init 2.91B 66.00 34.60 50.16 56.91 41.71 30.54 26.45 20.80 72.69 33.87 67.09 53.40 62.05 47.41 10.99
Pythia-12B Random Init 2.91B 67.00 31.80 27.32 50.36 38.18 27.40 25.25 21.21 67.85 27.65 57.79 40.58 60.15 41.73 13.26
Supernet Init 2.91B 69.00 33.20 41.76 51.46 40.79 29.29 26.20 21.21 70.57 28.67 61.99 45.98 58.04 44.47 11.71
Pythia-2.8B Random Init 2.78B 68.00 30.4 24.51 53.03 39.50 27.18 25.74 20.47 67.68 25.34 46.67 39.11 59.54 40.55 14.54
Table 9: Evaluation of Pythia models across multiple benchmarks. Reported numbers are metrics as defined in Section 4.5 (%).
Figures 1416 summarize the training behavior of the best models across different settings. The results highlight how architectures extracted from different search spaces (Figure 14), weight initialization strategies (Figure 15), and the use of distillation (Figure 16) affect convergence and final performance.
Teacher selection.
We also examine how the size of the teacher impacts student efficiency under the same compute budget (Figure 12). The smaller teacher model (Pythia-1B) achieves end-of-training perplexity comparable to, and marginally lower than, that of the larger Pythia-6.9B: $35.06$ vs. $35.15$ at 22B tokens. This indicates that, for our setup, alignment of inductive biases and gradient signal quality may be more critical than raw size of the teacher, and that excessively large teachers may be unnecessary under budget-constrained distillation.
Figure 12: Teacher size vs. student performance with a 2B-token budget. A Pythia-1B teacher achieves validation perplexity $35.06$, slightly better than Pythia-6.9B ($35.15$).
Figure 13: Evolutionary search on coarse uniform, coarse layer-wise, fine-grained uniform and fine-grained layer-wise search spaces for Pythia-6.9B. Minimum perplexity for each bin marked in red
Initialization #Params COPA OpenBookQA Lambada-OpenAI Winogrande Social IQA MMLU-cont. MMLU CommonsenseQA PIQA ARC-challenge ARC-easy HellaSwag BoolQ Avg-acc PPL-Nemotron-$c$
from-supernet 389M 61.00 30.00 24.02 51.54 38.23 26.21 26.35 18.84 65.34 24.57 51.38 33.11 60.73 39.33 23.66
from-supernet-distill 389M 66.00 30.60 23.95 49.57 37.41 26.31 25.62 19.57 65.56 25.34 49.66 33.92 54.31 39.06 18.59
from-supernet 1.04B 66.00 34.60 38.52 51.46 41.04 28.98 26.09 19.81 69.26 30.12 63.51 45.20 53.98 43.74 17.77
from-supernet-distill 1.04B 66.00 33.00 37.92 54.22 40.17 29.06 25.94 21.46 70.02 28.33 62.92 47.57 53.06 43.82 14.20
Table 10: Evaluation of sub-networks extracted from Pythia-6.9B for bin-1 and bin-2. Reported numbers are metrics as defined in Section 4.5 (%). We compare training with the cross entropy loss ( from-supernet) to training with knowledge distillation (from-supernet-distill) loss.
Figure 14: Training curves of the best models from each search space extracted from Pythia-6.9B (trained for 2 billion tokens)
Figure 15: Training curves of the best models found in each bin, initialized with supernet weights as well as random weights. A Pythia model of comparable size is also trained with random initialization in each bin as a baseline. The models are trained with 10 billion tokens.
Figure 16: Training curves of the best models in bins 1 and 2, obtained through distillation and pretraining. In both cases, the models are initialized with weights from the supernet (Pythia-6.9B). For comparison, a Pythia model of similar size is trained from random initialization in both bins, serving as a baseline. All models are trained on 10 billion tokens.
Algorithm 1 Bin-Constrained Evolutionary Search
1:Input: arch. space $\mathcal{S}$; bins $\{[L_{b},U_{b}]\}_{b=1}^{B}$; population NN; elites $k$; epochs $T$; random samples $r$; offspring $\lambda$; mutation prob. $m$; crossover prob. $c$
2:Helpers: $\mathrm{Params}(s)$ = param. count; $\mathsf{ppl}(s)$ = perplexity
3:Constrain(x,[L,U])(x,[L,U]): resample/repair until $\mathrm{Params}(x)\in[L,U]$
4:for $b=1$ to $B$ do
5:  $\mathcal{S}_{b}\leftarrow\{s\in\mathcal{S}\mid L_{b}\leq\mathrm{Params}(s)\leq U_{b}\}$
6:  Init population $\mathcal{P}^{(0)}_{b}\sim\mathcal{S}_{b}$
7:  for $t=0$ to T1T-1 do
8:   Evaluate $\mathsf{ppl}(s)$ for $s\in\mathcal{P}^{(t)}_{b}$
9:   Select elites $\mathcal{E}^{(t)}_{b}=\operatorname*{arg\,min}^{k}\mathsf{ppl}(s)$
10:   Mutants $\mathcal{O}_{\text{mut}}\leftarrow\textsc{Constrain}(\mathrm{Mut}(s),[L_{b},U_{b}])$, $s\sim\mathcal{E}^{(t)}_{b}$, size $\lambda$ (with prob. $m$)
11:   Crossovers $\mathcal{O}_{\text{cross}}\leftarrow\textsc{Constrain}(\mathrm{Cross}(s,s^{\prime}),[L_{b},U_{b}])$, $s,s^{\prime}\sim\mathcal{E}^{(t)}_{b}$, size $\lambda$ (with prob. $c$)
12:   Randoms $\mathcal{R}^{(t)}_{b}\sim\mathcal{S}_{b}$, size $r$
13:   Next pop. $\mathcal{P}^{(t+1)}_{b}\leftarrow\operatorname*{arg\,min}^{N}\mathsf{ppl}(s)$ over $\mathcal{E}^{(t)}_{b}\cup\mathcal{O}_{\text{mut}}\cup\mathcal{O}_{\text{cross}}\cup\mathcal{R}^{(t)}_{b}$
14:  end for
15:  Best $s_{b}^{\star}\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{P}^{(T)}_{b}}\mathsf{ppl}(s)$
16:end for
17:Output: $\{s_{b}^{\star}\}_{b=1}^{B}$
Appendix E Details on Importance Scoring
Importance scoring aims at defining scores for each transformer dimension, neuron or architecture parameter based on activation or weight magnitude. In our case, for a sub-network, the corresponding importance score serves as the proxy to sub-network quality or performance metrics like perplexity. Specifically, higher the importance score of a sub-network, superior the quality of the sub-network.
We adopt the dimension-wise importance scoring proposed by Muralidharan et al. [2024], which uses the activation of a component as proxy for its importance.
Given a batch as input ${\bm{X}}\in\mathbb{R}^{B\times T\times d_{model}}$ after applying the embedding layer ${\bm{W}}^{emb}$ we compute the following scores for each component, where $B$ corresponds to the batch dimension and $T$ corresponds to the sequence length dimension, and absabs corresponds to the absolute value function:
  • For a neuron $i\in\{1,...,U\}$ in a FFN layer $l$, we compute its importance by: $F^{(i)}_{FFN_{l}}=\nicefrac{{1}}{{B}}\sum_{B}\left(\nicefrac{{1}}{{T}}\sum_{T}\mathbf{X}{\bm{W}}_{1}^{l}[:,i]\right)$ where $\mathbf{W}_{1}^{l}[:,i]$ corresponds to all weights of neuron $i$ in layer $l$.
  • Similarly for each neuron $i\in\{1,...,d_{model}\}$ in the embedding layer we compute $F^{(i)}_{\text{emb}}=\nicefrac{{1}}{{B}}\sum_{B}\left(\nicefrac{{1}}{{T}}\sum_{T}\left(\text{Norm}(\mathbf{X}[:,:,i])\right)\right)$. Specifically we perform mean absolute aggregation over output of every (Layer or RMS) Norm layer as
  • For causal attention layers we compute the importance of head $h\in\{1,...,H\}$ of heads as :
    $F^{(h)}_{\text{MHA}}=\nicefrac{{1}}{{B}}\sum_{B}\Big(\nicefrac{{1}}{{T}}\sum_{T}\Big\|\text{Attn}\big({\bm{Q}}_{h},{\bm{K}}_{h},{\bm{V}}_{h}\big)\Big\|_{2}\Big)$
  • For a block $l\in\{1,...,L\}$ consisting of a MHA and a FFN layer with RMS or layer normalization in between, we compute the score: $\text{F}_{block}^{(l)}=1-\nicefrac{{1}}{{B}}\sum_{B}\Big(\nicefrac{{1}}{{T}}\sum_{T}\left(\frac{{\bm{X}}_{l}^{T}{\bm{X}}_{l+1}}{\|{\bm{X}}_{l}\|_{2}\|{\bm{X}}_{l+1}\|_{2}}\right)\Big)\;$ where ${\bm{X}}_{l}$ is the input to block $l$ and ${\bm{X}}_{l+1}$ the output.
Given, the score for each unit (layer, head or neuron), and a subnetwork configuration (for example: $e=64$, $d=128$, $l=2$, $h=4$), we compute the importance score corresponding the sub-network, by simply aggregating the normalized importance scores corresponding to the selected neurons, layers or heads. Similarly, for the weight space we define neuron, layer and head level importance scores by simply focussing on the magnitude of weight or neurons corresponding to every transformer dimension [Han et al., 2015].
Appendix F Whittle APIs
In addition to the functionalities core to our framework as defined in Section 3, we additionally define an api to compute different importance metrics for different sub-network dimensions. An overview of the Whittle library is shown in Figure 9.
compute_importance_score().
The compute_importance_score() function (Listing 6, Appendix F) assigns an importance score to a given sub-network, where higher scores indicate higher estimated quality. Importance scores are computed independently for each architectural component (e.g., layers, heads, neurons), normalized across available choices using a softmax, and aggregated by summation. This computation is performed once at the beginning of the search procedure, after which evaluating the importance of candidate sub-networks becomes inexpensive compared to full metrics such as perplexity.
Below we present the details of our API design for setting subnetwork 1, pretrain 2, convert to litgpt 4, distillation 5, 3 search and importance metric computation 6.
Listing 1: API for activating a sub-network
def set_sub_network(
sub_network_n_embd: int = 4, # Embedding dim
sub_network_intermediate_size: int | list[int] = 42 # MLP size
sub_network_num_heads: int | list[int] = 4, # No. of Attention heads
sub_network_n_layers: int = 4, # No. of Layers
sub_network_query_groups: int | list[int] = 2, # No. of Query groups
sub_network_head_size: int | list[int] = 4, # Head size
sampled_intermediate_indices: list[int] | list[list] = [4,8], # Sampled MLP neurons
sampled_head_indices: list[int] | list[list] = [2,3], # Sampled heads
sampled_query_group_indices: list[int] | list[list] = [0,1], # Sampled query groups
sampled_head_size_indices: list[int] | list[list] = [2,3,8,12], # Sampled head sizes
sampled_layer_indices: list[int] = [2,3,5,6], # Sampled layers
sampled_embd_indices: list[int] = [0,1,2,3], # Sampled embedding neurons
Listing 2: API for pretraining.
def pretrain(
model_name: str, # Name of the model to load. E.g., EleutherAI/pythia-410m
model_config: Optional[Config] = None, # Model configuration, overrides model_name
config_path : Optional[str] = None, # Path to yaml file with model configuration, overrides model_config
out_dir: Path = Path("out/pretrain"), # Path to save checkpoints to
precision: Literal["bf16-true", "bf16-mixed", "32-true", None] = None,
resume: Union[bool, Literal["auto"], Path] = False, # If true, resumes from the latest available checkpoint
data: Optional[DataModule] = None, # Dataset to use to train
train: TrainArgs = TrainArgs( # Training hyperparameters
save_interval=1000,
log_interval=1,
global_batch_size=512,
micro_batch_size=4,
max_tokens=int(3e12), # 3 trillion
max_norm=1.0,
min_lr=4e-5,
lr_warmup_steps=2000,
tie_embeddings=False,
eval: EvalArgs = EvalArgs(interval=1000, max_iters=100), # Evaluation hyper-parameters
optimizer: Union[str, Dict] = "AdamW", # Optimizer and its configuration
devices: Union[int, str] = "auto", # CUDA or CPU
num_nodes: int = 1, # Number of nodes for distributed training
tokenizer_dir: Optional[Path] = None, # Path to tokenizer (optional)
logger_name: Literal["wandb", "tensorboard", "csv", "mlflow"] = "tensorboard", # Logger to use
seed: int = 42, # Seed for reproducibility
init_from: str = "random", # Path to the checkpoint to load, or "random" to randomly initialize
use_flex: bool = False, # Set to True if the sub-network has layer-wise configuration
) -> LitGPT
Listing 3: API for supernetwork search.
def search(
supernet_name: str ="EleutherAI/pythia-12b", # litgpt model to extract SLM from
algorithm: str = "evolutionary_search", # name of search algorithm
num_bins: int = 4, # number of parameter bins
param_upper_bounds: list, # list of upper bounds for bins
param_lower_bounds: list, # list of lower bounds for bins
number_of_epochs: int, # number of epochs for search
)-> list[dict] # returns list of subnets
Listing 4: API for converting a subnet into a LitGPT model.
def convert_subnet_to_litgpt_model(
supernet_name: str = "EleutherAI/pythia-12b", # litgpt model to extract SLM from
subnet_config: dict = { # subnet configuration
sub_network_n_embd: 4,
sub_network_intermediate_size: 16,
sub_network_num_heads: 4,
sub_network_n_layers: 2,
sub_network_head_size: 4,
) -> LitGPT # returns LitGPT model
Listing 5: API for distilling a sub-network from a checkpoint.
def distill(
teacher_checkpoint_dir: Path, # Path to teacher model checkpoint directory
student_dir: Path, # Path to initialize student model directory with sub-network configuration and (optional) model weights
data: DataModule | None = None, # Dataset for distillation
out_dir: Path = Path("out/distill"), # Path to save distilled checkpoints
precision: Literal["bf16-true", "bf16-mixed", "32-true", None] = None, # Precision for training
train: TrainArgs = TrainArgs( # Training hyperparameters
save_interval=1000,
log_interval=1,
global_batch_size=512,
micro_batch_size=4,
max_tokens=int(5e8),
max_norm=1.0,
min_lr=4e-5,
lr_warmup_steps=2000,
tie_embeddings=False,
distill: DistillArgs = DistillArgs( # Distillation-specific hyperparameters
method="logits", # Distillation method (e.g., logits, hidden states)
temperature=10, # Softening factor for teacher logits
alpha=0.3, # Weight for student loss
beta=0.7, # Weight for distillation loss
loss="forward_kld", # Loss function for distillation
weight_scheme="other", # Weighting scheme for combining losses
eval: EvalArgs = EvalArgs(interval=50, max_iters=100, initial_validation=True), # Evaluation config
optimizer: str | dict = "AdamW", # Optimizer and configuration
devices: int | str = "auto", # CUDA or CPU
num_nodes: int = 1, # Number of nodes for distributed distillation
tokenizer_dir: Path | None = None, # Path to tokenizer (optional)
logger_name: Literal["wandb", "tensorboard", "csv"] = "csv", # Logger backend
seed: int = 42, # Seed for reproducibility
random_init_student: bool = False, # If True, randomly initialize student instead of loading
) -> LitGPT # Returns a distilled LitGPT student model
Listing 6: API for computing importance scores of subnet components.
def compute_importance_score(
supernet_name: str = "EleutherAI/pythia-12b", # base supernetwork
subnet_config: dict = { # subnet configuration
sub_network_n_embd: 4,
sub_network_intermediate_size: 16,
sub_network_num_heads: 4,
sub_network_n_layers: 2,
sub_network_head_size: 4,
layer_importance_type: str = "block importance", # method for layer scoring
head_importance_type: str = "minitron", # method for head scoring
neuron_importance_type: str = "minitron", # method for neuron scoring
) -> int: # returns importance score for a sampled sub-network
Appendix G Reproducibility Statements
We have taken extensive measures to ensure that all results in this paper can be replicated and verified by the community.
  • Code and Repository: We release all our code and scripts to reproduce our experiments at https://github.com/whittle-org/whittle/
  • Datasets and Pretrained Models: We evaluate on available benchmarks from lm-eval-harness (https://github.com/EleutherAI/lm-evaluation-harness) and use the publicly available Nemotron-CC dataset https://research.nvidia.com/labs/adlr/Nemotron-CC/ for training. Furthermore we use the Pythia-model suite, which is open-source https://github.com/EleutherAI/pythia.
  • Compute Resources: All our search experiments were run on on L40 GPU per parameter bin and base model. All our pretraining runs for bin-1 and bin-2 were run on 8 L40 GPUs and bin-3 was run on 4 H200 GPUs. All our distillation experiments were run on 4 H200 GPUs. We use cuda version 11.8.
  • Evaluation and Artifacts: Upon acceptance of the paper we will publicly release model checkpoints for all our experiments.